# Scaling Laws and Compute-Optimal Training Beyond Fixed Training Durations

### v2

@@ 6 Limitations
We conduct our experiments on models of up to 360M, training on up to 10B tokens. The trends we find are consistent across all scales, but the behavior may change at modern scales (Wei et al., 2022; Tay et al., 2021). Nonetheless, the approach of a constant learning rate followed by a cooldown has already proven successful at larger scales for released models (Hu et al., 2024; Shen et al., 2024; Zhai et al., 2022). Similarly, instabilities arising from a high learning rate for a long part of training can be alleviated (Wortsman et al., 2023). We therefore do not see any intrinsic barriers to the method at scale.
Secondly, we exclusively focus on the training or validation loss, though downstream tasks are ultimately the main metric of interest. While they often scale reliably with the loss (Du et al., 2024; Gadre et al., 2024), this remains an important question for further research; in particular, when coupled with different data-mixtures during the cooldown phase, as is the focus of (Hu et al., 2024; Shen et al., 2024) and out of the scope of our work.
@@ 7 Related work
Cosine Schedules and Alternatives for Transformers. The cosine decay was originally introduced by Loshchilov & Hutter (2016) for cyclic schedules in vision tasks, where it is common practice to have stepwise or cyclic LRs to escape bad minima when training multiple epochs (Smith et al., 2017). For language models where data is more vast, the cosine schedule with a single cycle is currently the de-facto standard for training, with few exceptions of T5 (Raffel et al., 2020) and PaLM1 (Chowdhery et al., 2023) that used a form of inverse square root. In line with our work, recently released models opt for alternatives such as stepwise schedules (Bi et al., 2024) or the presented constant + cooldown (Shen et al., 2024; Hu et al., 2024). These alternatives were previously also explored for vision transformers by Zhai et al. (2022), who find reciprocal square-root with cooldown to perform best, and in the context of continual learning (Ibrahim et al., 2024; Gupta et al., 2023). Defazio et al. (2023) investigate the gap between theory and practice of learning rate schedules and suggest that a linear decay is optimal, also for LLM training. We similarly find that a long linear decay can slightly outperform cosine.
@@
Weight Averaging. Weight averaging over past iterates (Polyak & Juditsky, 1992) has long been known to be beneficial for convergence. Izmailov et al. (2018) introduce stochastic weight averaging for better generalization in deep learning models. Similarly, exponential moving average is commonly used in vision (Morales-Brotons et al., 2024). Importantly, Sandler et al. (2023) show equivalency of WA to decaying learning rate schedules. In the context of LLM training, close to our work is Sanyal et al. (2023) which showed that a form of latest averaging (Kaddour, 2022) can be used to improve the performance of models early in training. However, they did not investigate the relation of weight averaging to compute optimality and its implications for scaling experiments.
@@
Scaling Law Experiments for Neural Language Models. Kaplan et al. (2020) were the first to establish scaling laws for language models by training a suite of models for a fixed token count. Important to our work, Hoffmann et al. (2022) revise these findings and demonstrate specific methods to establish scaling laws, notably training a family of models for different cosine lengths. The subsequent models like LLama and LLama2 (Touvron et al., 2023a, b) further improve performance of smaller models by training beyond the Chinchilla optimal point, motivated by lower inference costs (Gadre et al., 2024; De Vries, 2023; Sardana & Frankle, 2023). Recent works (Muennighoff et al., 2023; Bi et al., 2024; Goyal et al., 2024) highlight how data repetition and quality affect the scaling behavior, which suggests that scaling laws should be updated more frequently. However, these works do not consider efficient experiments for scaling laws, which is the focus of our work.
@@ 8 Conclusion
We believe the results are of great importance to the present and future of LLM training: the presented methods facilitate research for the current post-Chinchilla era, where models are trained much beyond compute-optimal, by allowing more flexibility to continue training whenever needed. At the same time, recent results that suggest data-dependency in scaling (Bi et al., 2024; Goyal et al., 2024; Aghajanyan et al., 2023; Pandey, 2024) imply the need to frequently update scaling laws, which is economically more feasible with reduced costs. We therefore hope that our work will make scaling research more accessible to researchers and practitioners alike.
@@ Acknowledgements
We thank Anastasia Koloskova, Amirkeivan Mohtashami and Thomas Wolf for helpful discussions regarding the paper and its implications.
@@ A.1 Overview
Implementation and Infrastructure. Our code is based on an extension of NanoGPT [^2] https://github.com/karpathy/nanoGPT and uses PyTorch (Paszke et al., 2017) as well as FlashAttention (Dao et al., 2022). We incorporate the bfloat16 for memory and throughput, trained with mixed precision float32 parameters and bfloat16 activations (Micikevicius et al., 2017). All our experiments were performed using a cluster of A100 GPUs (both 40GB/80GB RAM) with 2 data-parallel (i.e., 2 GPUs per run). A few selected runs used a single node of 8 H100s. We estimate that the full cost of all experiments for this project (including prototyping) amounts to roughly 2500 GPU hours.
@@
Table 1: Model configurations. We provide an overview of the model sizes and hyperparameters for the different models in the scaling experiments.
Model LR (Cos/Const) Batch Size Steps Tokens Token/Params Ratio
@@
A.2 FLOPS computation
@@ B.1 More Results on Cooldown
Length of cooldown. We repeat our ablation from Sect. 3 in Figure 16 with the absolute number of decay steps (not fractional) in Figure 17.
@@
Figure 15: The order and results of different cooldowns hold across settings. Left: We change the learning rate to $5e^{-4}$ instead of $1e^{-3}$ as in previous experiments. Right: The performance for 10% cooldown steps instead of 20%.
Figure 16: Parabola shape of the relationship between cooldown length and final perplexity. We repeat the plots from Fig. 5 with a zoomed-in view on the left.
@@
B.2 Learning Curves for All Models of Scaling Experiments
@@
B.3 Additional Results and Compute Savings.
Figure 21: FLOPS Savings for all Models in Scaling Experiments across different Runs (Log Scale). The FLOPS savings amount to a factor of $\frac{1}{2}$ across all models. If more runs across different lengths were to be performed (e.g., 3 per model in our experiments), the difference becomes more significant.
@@ B.4 Additional Experiments on OpenWebText2
Figure 23: Results Transfer to OpenWebText2. We see the same behavior for cooldown schedules and SWA, verifying the reliability of the alternative training methods.

### v3

@@ 6 Limitations
We conduct our experiments on models of up to 8B parameters, with long training runs of a 1B model on multiple hundred tokens. The trends we find are consistent across all scales, but training behavior can be more brittle at modern scales and extremely long training (Wei et al., 2022; Tay et al., 2021). However, instabilities arising from a high learning rate for a large part of training can be alleviated (Wortsman et al., 2024).
@@ 7 Related work
Cosine Schedules and Alternatives for Transformers. The cosine decay was originally introduced by Loshchilov & Hutter (2016) for cyclic schedules in vision tasks, where it is common practice to have stepwise or cyclic LRs to escape bad minima when training multiple epochs (Smith et al., 2017). For language models where data is more vast, the cosine schedule with a single cycle is currently the de-facto standard for training, with few exceptions of T5 (Raffel et al., 2020) and PaLM1 (Chowdhery et al., 2023) that used a form of inverse square root. In line with our work, recently released models opt for alternatives such as stepwise schedules (Bi et al., 2024) or the presented constant + cooldown (Shen et al., 2024; Hu et al., 2024). These alternatives were previously also explored for vision transformers by Zhai et al. (2022), who find reciprocal square-root with cooldown to perform best, and in the context of continual learning (Ibrahim et al., 2024; Gupta et al., 2023). Defazio et al. (2023) investigate the gap between theory and practice of LR schedules and suggest that a linear decay is optimal, also for LLM training. We similarly find that a long linear decay can slightly outperform cosine.
@@
Weight Averaging. Weight averaging over past iterates (Polyak & Juditsky, 1992) has long been known to be beneficial for convergence. Izmailov et al. (2018) introduce stochastic weight averaging for better generalization in deep learning models. Similarly, the exponential moving average is commonly used in vision (Morales-Brotons et al., 2024). Importantly, Sandler et al. (2023) show the equivalence of WA with decaying learning rate schedules. In the context of LLM training, close to our work is Sanyal et al. (2023) which showed that a form of latest averaging (Kaddour, 2022) can be used to improve the performance of models early in training. However, they did not investigate the relation of weight averaging to compute optimality and its implications for scaling experiments.
@@
Scaling Law Experiments for Neural Language Models. Kaplan et al. (2020) were the first to establish scaling laws for language models. Important to our work, Hoffmann et al. (2022) revise these findings and demonstrate specific methods to establish scaling laws, notably training a family of models for different cosine lengths. The subsequent models like LLama1/2 (Touvron et al., 2023a, b) further improve performance of smaller models by training beyond the Chinchilla optimal point, motivated by lower inference costs (Gadre et al., 2024; De Vries, 2023; Sardana & Frankle, 2023). Recent works (Muennighoff et al., 2023; Bi et al., 2024; Goyal et al., 2024) highlight how data repetition and quality affect the scaling behavior, which suggests that scaling laws should be updated more frequently. However, these works do not consider efficient experiments for scaling laws, which is the focus of our work.
In concurrent work, Porian et al. (2024) and Pearce & Song (2024) find that the discrepancy between the Kaplan and Chinchilla scaling laws is not attributed to the learning rate decay, but they establish that the optimal token/parameter ratio can be obtained with a constant learning rate without any cooldown; however, the actual performance is then suboptimal, and a cooldown schedule as suggested in our work is needed to properly estimate model performance, in particular for downstream tasks.
@@ 8 Conclusion
We believe the results are of great importance to the present and future of LLM training: The presented methods facilitate research for the current post-Chinchilla era, where models are trained much beyond compute-optimal, by allowing more flexibility to continue training whenever needed. At the same time, recent results that suggest data dependency in scaling (Bi et al., 2024; Goyal et al., 2024; Aghajanyan et al., 2023; Pandey, 2024) imply the need to frequently update scaling laws, which is economically more feasible with reduced costs. We therefore hope that our work will make scaling research more accessible to researchers and practitioners alike.
@@ Acknowledgements
We thank Anastasia Koloskova, Amirkeivan Mohtashami and Thomas Wolf for helpful discussions regarding the paper and its implications. We also thank Stella Biderman, whose open discussions [^2] https://x.com/BlancheMinerva/status/1754559681933324372 inspired parts of our experiments.
@@ A.1 Overview
Implementation and infrastructure. Our code is based on an extension of NanoGPT [^3] https://github.com/karpathy/nanoGPT and uses PyTorch (Paszke et al., 2017) as well as FlashAttention (Dao et al., 2022). We incorporate bfloat16 for memory and throughput, trained with mixed precision float32 parameters and bfloat16 activations (Micikevicius et al., 2017). All experiments (aside from 1B and 8B, see A.2) were performed using a cluster of A100 GPUs (both 40GB/80GB RAM) with 2 data-parallel (i.e. 2 GPUs per run). Some selected runs used a single node of 8 H100s. We estimate that the full cost of all experiments for this project (including prototyping) to amount to roughly 2500-3000 GPU hours.
@@
Table 2: Model configurations for scaling law experiments. We provide an overview of the model sizes and hyperparameters for the different models in the scaling experiments.
A.2 Setup for the 1B and 8B Runs
Setup. In order to efficiently scale to larger models, we use the nanotron library  [^4] https://github.com/huggingface/nanotron that enables all forms of training parallelism while being performant and flexible. The configuration and hyperparameters of the two models are described in Table 4. Similarly to our other experiments, we use the most common setup of the AdamW optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$, weight decay of $0.1$, and gradient clipping with $1.0$. For the 1B model, we use a custom BPE tokenizer [^5] https://hf.co/lvwerra/the-tokenizer-v1 (that is very similar to GPT-2) because the FineWeb dataset (Penedo et al., 2024) was already pre-tokenized on our cluster. Each run for the 1B model was performed on 4xH100s GPUs. For the 8B model, we reuse the original Llama3 tokenizer (Dubey et al., 2024) for the similar reason that FineWeb-Edu was pre-tokenized on our cluster; here, we use 12 nodes, each composed of 4xGH200 GPUs. The model is split within one node with tensor parallel (TP) set to 4.
Evaluation. For evaluation, we use the lighteval library (Fourrier et al., 2023) that works directly with nanotron checkpoints. We run the most common LLM benchmarks including MMLU (Hendrycks et al., 2021), ARC (Clark et al., 2018), OpenBookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019), CommonSenseQA (Talmor et al., 2019), SIQA (Sap et al., 2019), Winogrande (Sakaguchi et al., 2021), truncating large benchmarks to 1000 samples so that we could efficiently evaluate over the course of training. This is a custom setup that is equal to that used for FineWeb ablations; the setup to reproduce the eval is described here [^6] https://github.com/huggingface/cosmopedia/blob/d62abb8e13c567b999e33d0a1d795968bc052c6a/evaluation/README.md  [^7] https://hf.co/datasets/HuggingFaceFW/fineweb/blob/main/lighteval_tasks.py#L12. We report the accuracy normalized by sequence length (acc_norm).
Model LR (Cos/Const) BS Steps Tokens Token/Params Ratio
@@ A.2 Setup for the 1B and 8B Runs
Parameter 1B 8B
d_model 1792 4096
n_layers 24 32
ffw_size 4864 14336
kv_size 128 128
n_heads 14 32
n_kv_heads 14 8
vocab_size 49152 128256
seq_len 2048 4096
warmup_steps 2000 1000
batch_size (1.8M, 2M) 0.6M
Total Steps (55k, 220k) 20k
[…]
Table 4: Model configurations for larger runs. The batch size and learning rate for the 1B model tokens were estimated using DeepSeek scaling laws for 100B tokens. The two values for BS and the total steps of the 1B model distinguish the runs (100B,460B). For the 8B model, the architecture is identical to Llama3 and the batch size was set according to the available GPU limits on our cluster at the time of running experiments.
A.3 FLOPs computation
@@ B.1 More Results on Cooldown
Note that the order might change for substantially different cooldown lengths; we focus on 10% and 20% because they are practically relevant.
In addition, we perform a comparison of different exponents for the square-root function: since the (1-Sqrt) can be expressed as $(1-x^{a})$ where $a=0.5$, we sweep the parameter $a<0.5$. The results are shown in Figure 18. Apart from 0.1 and 0.2, which perform noticeably worse because the learning rate is too low for many steps, the other exponents only show a marginal difference to $a=0.5$, which still comes out on top.
We aim to further investigate the impact of the functional form of the cooldown in future work.
Length of cooldown. We repeat our ablation from Sect. 3 with a 210M model in Figure 19 (fractional x-axis) and with the absolute number of decay steps (not fractional) in Figure 20.
@@
Figure 17: The order and results of different cooldowns hold across settings. Left: We change the learning rate to $5e^{-4}$ instead of $1e^{-3}$ as in previous experiments. Right: The performance for 10% cooldown steps instead of 20%. Note that the order might change for substantially different cooldown lengths; we focus on 10% and 20% as they are practically relevant.
Figure 18: Different exponents for functions. Since the negative square root performs remarkably well, we experiment with varying the exponent of the functional form $(1-x^{a})$ for $a<0.5$. Besides 0.1 and 0.2 which perform noticeably worse, the other exponents only show marginal difference in this experiment, with $a=0.5$ (the square root) still coming out on top.
Figure 19: Parabola shape of the relationship between cooldown length and final perplexity. We repeat the experiment from Fig. 5 with a 210M parameter model and the zoomed-in view on the left.
@@
Annealing cosine to less than 10%. We ablate the choice of the final learning rate for the cosine annealing in Fig. 22, where we find that the final value should be set lower than simply 10% of the maximum. However, when evaluating on downstream benchmarks (see Section B.5), we see that annealing to zero can hurt metrics; we posit this comes from too early saturation. Combining these two arguments implies that the maximum and final LR should be set (and ideally swept over) independently.
Figure 22: The cosine schedule should be decayed to lower values. While setting the minimum to 10% of the maximum LR is the most common heuristic, the final LR should be set lower and independently of the maximum. In our experiments, cosine then matches the cooldown with (1-Sqrt) (left; optimal LR shown), albeit the LR sensitivity remains (right; larger maximum LR led to divergence).
B.2 Additional Results and Compute Savings
Figure 24: FLOPs savings for all models in scaling experiments across different runs (log scale). The savings amount to a factor of $\frac{1}{2}$ across all models. If more runs across different lengths were to be performed (compared to 3 per model in our experiments), the difference would be even more significant.
B.3 Learning Curves for All Models of Scaling Experiments
@@ B.4 Additional Experiments on OpenWebText2
Figure 27: Results Transfer to OpenWebText2. We see the same behavior for cooldown schedules and SWA, verifying the reliability of our findings.
B.5 Full Results of Large Model Runs
In this section, we report the detailed results of the 1B runs with the downstream benchmarks.
In Figure 28, we compare the aggregate metrics throughout training for both the 100B (left) and the 460B token run (right). We equally plot individual benchmark curves in Figure 29 for the 100B run with a zoomed view after 80B tokens in Figure 30; interestingly, we observe a similar uptick in performance for some metrics (e.g., MMLU, HellaSwag) with the cooldown, while others do not benefit as clearly (e.g., OpenBookQA). This observation is an interesting direction for further research.
The final numbers are given in Table 5 (100B) and Table 6 (460B). Notably, longer cooldowns do not necessarily improve the metrics (e.g. going from 5% to 20%).
Figure 28: Aggregate metrics throughout training of the 1B model on 100B and 460B tokens. We train a 1B model on 100B (left) and 460B tokens (right) of FineWeb (Penedo et al., 2024), and find that the performance of cosine and cooldown matches. Though cosine to zero improves the loss (Figure 22), it leads to a saturation before the end of training, hurting overall performance.
Figure 29: Detailed benchmarks throughout training of the 1B model on 100B tokens. For the cooldown (starting at 80B tokens), we observe a similar uptick in performance for some metrics (e.g., MMLU, HellaSwag) akin to the observed drop in loss. Other metrics do not benefit as clearly from the cooldown (e.g., OpenBookQA).
Figure 30: Zoomed-in view of the 1B model benchmarks after 80B tokens. We repeat Figure 29 with a focus on the cooldown phase, which starts at 80B tokens.
Metric Cosine to 10% Cosine to 0 1-Sqrt 20% Linear 20%
Aggregated Score 46.26 45.88 46.23 46.20
MMLU 30.21 30.70 30.84 30.71
ARC 41.25 42.35 41.80 41.25
OpenBookQA 35.40 33.80 36.00 35.80
PIQA 76.10 75.50 76.50 77.00
HellaSwag 56.30 55.60 55.80 55.90
CommonSenseQA 36.10 36.20 37.00 36.70
SIQA 40.70 40.20 39.60 39.50
Winogrande 54.00 52.70 52.30 52.70
Table 5: Final evaluation results after 100B tokens. Both cosine and the cooldown schedules have comparable final numbers, with only slight differences for certain benchmarks.
Metric Cosine to 0 1-Sqrt 5% Linear 5% Linear 10% Linear 20%
Aggregated Score 48.03 47.91 47.84 47.98 47.92
MMLU 31.25 31.71 31.65 31.78 31.84
ARC 45.20 44.10 44.05 44.15 45.05
OpenBookQA 37.60 37.80 38.00 38.20 37.20
PIQA 78.10 77.40 77.40 77.30 77.80
HellaSwag 59.90 59.60 59.50 60.00 59.20
CommonSenseQA 37.70 37.30 36.70 36.90 36.90
SIQA 39.90 39.50 39.40 39.50 39.70
Winogrande 54.60 55.90 56.00 56.00 55.70
Table 6: Final evaluation results after 460B tokens. The findings of Table 5 transfer to much longer training runs with 460B tokens, where the performances of cosine and cooldowns match well. Notably, longer cooldowns do not necessarily improve the metrics (e.g. going from 5% to 20%).